OpenStudy (anonymous):

Which is larger? 1 millimeter 1 centimeter 1 kilometer 1 decameter

OpenStudy (anonymous):

1 kilometer

OpenStudy (xishem):

Well, the prefixes mean... Milli = 1*10^(-3) Centi = 1*10^(-2) Kilo = 1*10^(3) Deca = 1*10^(1) Which of those is the largest?

OpenStudy (anonymous):

kilo